OpenStudy (anonymous):

Could someone help me with Geometry? I would be super grateful. :) What is the volume of a square pyramid with base edges of 20 meters and a height of 30 meters? A. 1,800 m³ B. 4,000 m³ C. 600 m³ D.12,000 m³

jimthompson5910 (jim_thompson5910):

V = (Area of base)*(Height)/3 V = (20^2)*(30)/3 V =400*10 V = 4000 So the volume is 4,000 cubic meters

OpenStudy (apoorvk):

the volume of any pyramid is V= (1/3)*(area of base)*(height) now can you solve the problem??

OpenStudy (anonymous):

You people are absolutely incredible!! Thank you so much! You both got very deserving medals ;)

OpenStudy (anonymous):

If anyone else who was going to answer this question wants to write anything, anything at all- I'll give you a medal because you took the time to at least look at it. And for that I thank you very much :)

OpenStudy (apoorvk):

heyy mate. don't thank us. it's what we are all here for, innit? :)